# ClimaQA: An Automated Evaluation Framework for Climate Foundation Models

### v1

@@ 5 Experiments
5.1 Experimental Setup
We evaluate different families of LLMs on our benchmark. We use TogetherAI for performing inference on open source models like gemma-27b (Team et al., 2024), llama3-70b(Dubey et al., 2024), and mixtral-8x22b (Jiang et al., 2024). We also evaluate OpenAI’s (Achiam et al., 2023) gpt-3.5-turbo and gpt-4o.
We evaluate each of these models in 3 settings - default, few-shot prompting (Brown, 2020), and Retrieval Augmented Generation (Lewis et al., 2020). For the MCQs, the models were prompted to output a single letter representing the correct option, and the top-most token was chosen as the answer. For Freeform QA, the models were prompted to output concise answers with a maximum of 2 sentences. For Cloze QA, the models were prompted to output a single scientific word that best fits the blank with respect to the context around it.
Additionally, we conduct further pre-training on graduate-level textbook data for both the LLaMA3.1-8B-Instruct and Mistral-7B-v0.3-Instruct (Jiang et al., 2023) models. This continued pre-training was based on 13 distinct graduate textbooks that were not part of the question-generation process. The objective was to enhance the model’s climate knowledge without directly exposing it to the specific sources used for question generation, thereby reducing the risk of data contamination.
Finally, we fine-tune LLaMA3.1-8B-Instruct and Mistral-7B-v0.3-Instruct on the ClimaQA-Silver dataset, which contains all three forms of MCQ, Freeform, and Cloze, in different complexity levels. We then evaluate the impact of this task-specific fine-tuning by assessing the models’ performance on the ClimaQA-Gold dataset.
The details of the continued pre-training and fine-tuning procedure are explained in Appendix A.2.
5.2 Results
We report the performance of various models across different QA forms and complexities. Table 3 shows the results for the MCQ and Cloze form questions, and free-form results are demonstrated in Table 4, in Appendix A.1. We observe that most models struggle with reasoning MCQs compared to base and hypothetical questions. While some models perform poorly on reasoning and hypothetical MCQs, they tend to generate stronger responses for the same type of Freeform Questions, indicating improved performance when reasoning is emphasized over factual recall similar to observations of Chain of Thought experiments (Wei et al., 2022), as shown in Figure 5.
@@ 5.2 Results
Additionally, we examine the impact of providing relevant context through RAG when answering questions. We explore two retrieval scenarios: one where the model retrieves from 13 books that were not used to generate the questions (rag-13), and another where it retrieves from the five books that the questions were derived from (rag-5). Retrieval from the five textbooks consistently improves performance across all tasks, whereas retrieval from the 13 books shows less predictable results due to the potential presence of distracting contexts. Furthermore, Few-shot prompting yields marginal improvements in most cases.
@@
Continued pretraining on the graduate textbooks leads to improved performance in both MCQs and Cloze question-answering tasks. Additionally, fine-tuning on the ClimaQA-Silver dataset further enhances performance, often producing the best results after RAG on the five textbooks (rag-5).
@@
We observe that the BLEU and BERTScore metrics are slightly biased towards the model that was used for QA-generation (gpt-3.5-turbo) while this is not seen in the proposed Factual Accuracy metric 5. Overall, GPT-4o dominates across tasks, demonstrating superior performance compared to other models in this evaluation set.
Table 3: Performance analysis of various state of the art LLMs on MCQs and Cloze QA. Bold marks the max within a model’s variants and green highlights the overall column max.
Model MCQ Cloze
Base Reason Hypo Overall EM PS
gemma-27b 82.02 66.67 84.21 77.90 49.38 0.87
gemma-27b-fs 80.90 74.07 84.21 79.56 52.50 0.88
gemma-27b-rag-5 87.64 77.78 81.58 83.43 56.88 0.90
gemma-27b-rag-13 80.90 72.22 84.21 79.01 45.62 0.85
gpt-3.5-turbo 78.65 64.81 73.68 73.48 43.12 0.81
gpt-3.5-turbo-fs 79.78 74.07 76.32 77.35 43.75 0.78
gpt-3.5-turbo-rag-5 77.53 81.48 78.95 79.01 68.75 0.92
gpt-3.5-turbo-rag-13 66.29 68.52 71.05 67.96 39.38 0.81
gpt-4o 84.27 81.48 84.21 83.43 53.12 0.88
gpt-4o-fs 83.15 83.33 84.21 83.43 56.25 0.89
gpt-4o-rag-5 94.38 87.04 92.11 91.71 71.88 0.94
gpt-4o-rag-13 78.65 83.33 84.21 81.22 50.62 0.88
llama3-70b 80.90 75.93 81.58 79.56 38.75 0.82
llama3-70b-fs 79.78 75.93 81.58 79.01 48.12 0.85
llama3-70b-rag-5 92.13 81.48 86.84 87.85 63.12 0.91
llama3-70b-rag-13 78.65 72.22 84.21 77.9 43.75 0.84
mixtral-8x22b 78.16 74.07 81.58 77.65 35.62 0.75
mixtral-8x22b-fs 78.16 77.78 78.95 78.21 45.00 0.83
mixtral-8x22b-rag-5 89.66 75.93 81.58 83.80 45.00 0.78
mixtral-8x22b-rag-13 79.31 66.67 68.42 73.18 28.12 0.65
mistral-7b 66.29 59.26 81.58 67.4 17.50 0.74
mistral-7b-fs 66.29 61.11 71.05 65.75 33.12 0.80
mistral-7b-rag-5 92.13 81.48 81.58 86.74 57.5 0.88
mistral-7b-rag-13 71.91 62.96 73.68 69.61 22.50 0.76
mistral-7b-book 70.79 64.81 76.32 70.17 21.25 0.76
mistral-7b-qa 74.16 70.37 76.32 73.48 41.88 0.83
llama3.1-8b 69.66 55.56 68.42 65.19 26.25 0.77
llama3.1-8b-fs 75.28 68.52 76.32 73.48 38.75 0.82
llama3.1-8b-rag-5 93.26 77.78 89.47 87.85 72.5 0.92
llama3.1-8b-rag-13 76.4 64.81 76.32 72.93 36.25 0.81
llama3.1-8b-book 75.28 70.37 73.68 73.48 30.00 0.77
llama3.1-8b-qa 79.78 64.81 78.95 75.14 51.25 0.85
@@ 6 Conclusion
The ClimaQA benchmark offers a comprehensive framework for evaluating foundation models in climate question-answering, addressing critical aspects such as reasoning, factual accuracy, and understanding of scientific terminology. By incorporating freeform, multiple-choice, and cloze task forms with different levels of complexity, the benchmark rigorously tests models across different dimensions of scientific inquiry. Furthermore, the use of advanced metrics, such as factual accuracy for freeform tasks and phrase similarity for cloze tasks, can provide a more nuanced assessment of model performance.
@@ Appendix A Appendix
A.1 Freeform Evaluation Results
Table 4: Performance analysis of various state of the art LLMs on Freeform QA. Bold marks the max within a model’s variants and green highlights the overall column max.
@@ B R H O B R H O B R H O
gemma-27b-rag-5 0.462 0.573 0.552 0.529 0.902 0.911 0.906 0.906 0.84 0.85 0.85 0.85
gemma-27b-rag-13 0.349 0.496 0.524 0.456 0.867 0.889 0.896 0.885 0.58 0.73 0.77 0.69
@@
gpt-3.5-turbo-rag-5 0.528 0.601 0.561 0.563 0.908 0.919 0.921 0.916 0.85 0.86 0.87 0.86
gpt-3.5-turbo-rag-13 0.425 0.477 0.543 0.482 0.877 0.889 0.904 0.891 0.46 0.64 0.81 0.64
@@
llama3-70b-rag-5 0.441 0.574 0.604 0.540 0.909 0.916 0.917 0.914 0.82 0.81 0.84 0.82
llama3-70b-rag-13 0.341 0.475 0.519 0.445 0.872 0.890 0.902 0.888 0.45 0.62 0.79 0.62
@@
mixtral-8x22b-rag-5 0.458 0.598 0.610 0.555 0.905 0.916 0.912 0.911 0.78 0.82 0.86 0.82
mixtral-8x22b-rag-13 0.341 0.488 0.546 0.459 0.870 0.890 0.898 0.886 0.50 0.62 0.77 0.63
@@
llama3.1-8b-rag-5 0.509 0.550 0.534 0.531 0.905 0.920 0.911 0.912 0.77 0.80 0.83 0.80
llama3.1-8b-rag-13 0.392 0.469 0.507 0.456 0.873 0.893 0.901 0.889 0.52 0.65 0.72 0.63
llama3.1-8b-book 0.420 0.455 0.496 0.457 0.876 0.892 0.902 0.890 0.53 0.70 0.71 0.65
llama3.1-8b-qa 0.436 0.469 0.533 0.480 0.882 0.896 0.904 0.894 0.53 0.59 0.77 0.63
@@
mistral-7b-rag-5 0.541 0.455 0.448 0.482 0.898 0.912 0.907 0.906 0.77 0.82 0.84 0.81
mistral-7b-rag-13 0.413 0.356 0.399 0.390 0.868 0.884 0.896 0.883 0.54 0.60 0.73 0.62
mistral-7b-book 0.295 0.176 0.189 0.221 0.880 0.890 0.899 0.890 0.52 0.70 0.70 0.64
mistral-7b-qa 0.520 0.386 0.434 0.447 0.892 0.900 0.906 0.899 0.40 0.61 0.82 0.61
A.2 Training Details

### v2

@@ 5 Experiments
5.1 Experimental Setup
We evaluate different families of LLMs on our benchmark. We use TogetherAI for performing inference on open source models like gemma-27b (Team et al., 2024b), llama3-70b(Dubey et al., 2024), and mixtral-8x22b (Jiang et al., 2024). We also evaluate OpenAI’s (Achiam et al., 2023) gpt-3.5-turbo and gpt-4o.
We evaluate each of these models in 3 settings - default, few-shot prompting (FS)(Brown, 2020), and Retrieval Augmented Generation (RAG) (Lewis et al., 2020). For the MCQs, the models were prompted to output a single letter representing the correct option, and the top-most token was chosen as the answer. For Freeform QA, the models were prompted to output concise answers with a maximum of 2 sentences. For Cloze QA, the models were prompted to output a single scientific word that best fits the blank with respect to the context around it.
We conduct further pre-training on graduate-level textbook data for both the LLaMA3.1-8B-Instruct and Mistral-7B-v0.3-Instruct (Jiang et al., 2023) models. This pre-training was based on 13 distinct graduate textbooks that were not part of the question-generation process. The objective was to enhance the model’s climate knowledge without directly exposing it to the specific sources used for question generation, thereby reducing the risk of data contamination.
Table 3: Performance analysis of various state-of-the-art LLMs on MCQs and Cloze QA. While source represents the set of books used for QA generation, held-out represents the remaining set of books. Bold marks the max within a model’s variants and green highlights the overall column max.
Model MCQ Cloze
Base Reason Hypo Overall EM PS
gemma-27b 81.75 72.22 82.98 79.18 49.38 0.87
gemma-27b-fs 80.95 77.78 82.98 80.41 52.50 0.88
gemma-27b-rag-source 90.48 80.56 78.72 85.31 56.88 0.90
gemma-27b-rag-held-out 79.37 76.39 78.72 78.37 45.62 0.85
gpt-3.5-turbo 74.34 69.91 74.47 73.06 43.12 0.81
gpt-3.5-turbo-fs 76.98 74.54 76.60 76.19 43.75 0.78
gpt-3.5-turbo-rag-source 80.42 80.09 77.30 79.73 68.75 0.92
gpt-3.5-turbo-rag-held-out 70.63 71.30 69.50 70.61 39.38 0.81
gpt-4o 86.77 86.11 82.27 85.71 53.12 0.88
gpt-4o-fs 87.83 87.50 80.85 86.39 56.25 0.89
gpt-4o-rag-source 95.77 91.67 86.52 92.79 71.88 0.94
gpt-4o-rag-held-out 82.80 80.56 81.56 81.90 50.62 0.88
llama3-70b 84.92 80.56 82.98 83.27 38.75 0.82
llama3-70b-fs 82.54 81.94 82.98 82.45 48.12 0.85
llama3-70b-rag-source 92.06 84.72 87.23 88.98 63.12 0.91
llama3-70b-rag-held-out 80.95 76.39 85.11 80.41 43.75 0.84
mixtral-8x22b 80.16 79.17 80.85 80.00 35.62 0.75
mixtral-8x22b-fs 80.95 81.94 80.85 81.22 45.00 0.83
mixtral-8x22b-rag-source 90.48 80.56 76.60 84.90 45.00 0.78
mixtral-8x22b-rag-held-out 80.16 73.61 74.47 77.14 28.12 0.65
mistral-7b 64.29 63.89 82.98 67.76 17.50 0.74
mistral-7b-fs 65.08 66.67 76.6 67.76 33.12 0.80
mistral-7b-rag-source 92.86 84.72 85.11 88.98 57.5 0.88
mistral-7b-rag-held-out 66.67 62.5 74.47 66.94 22.50 0.76
mistral-7b-cp-held-out 72.22 62.5 74.47 69.8 21.25 0.76
mistral-7b-ft-silver 73.81 75.00 80.85 75.51 41.88 0.83
llama3.1-8b 76.98 62.50 65.96 70.61 26.25 0.77
llama3.1-8b-fs 76.19 72.22 76.60 75.10 38.75 0.82
llama3.1-8b-rag-source 94.44 83.33 89.36 90.2 72.50 0.92
llama3.1-8b-rag-held-out 76.19 66.67 76.6 73.47 36.25 0.81
llama3.1-8b-cp-held-out 77.78 75.00 72.34 75.92 30.00 0.77
llama3.1-8b-ft-silver 74.6 70.83 72.34 73.06 51.25 0.85
gemini-1.5-flash 82.54 73.61 87.23 80.82 50.62 0.88
gemini-1.5-flash-long-cxt-source 88.1 75.00 80.85 82.86 51.88 0.89
gemini-1.5-flash-long-cxt-held-out 70.63 70.83 78.72 72.24 46.88 0.87
Additionally, we fine-tune LLaMA3.1-8B-Instruct and Mistral-7B-v0.3-Instruct on the ClimaQA-Silver dataset, which contains all three forms of MCQ, Freeform, and Cloze, in different complexity levels. We then evaluate the impact of this task-specific fine-tuning by assessing the models’ performance on the ClimaQA-Gold dataset. The details of the continued pre-training and fine-tuning procedure are explained in Appendix A.3. We also leverage Gemini 1.5 (Team et al., 2024a), with a context window of up to 1 million tokens, to pass an entire textbook in context and answer questions based on that. Finally, we evaluate models on the ClimaQA-Silver dataset and analyze its potential differences from ClimaQA-Gold in Appendix A.2.2.
5.2 Results
We report the performance of various models across different QA forms and complexities. Table 3 shows the results for the MCQ and Cloze form questions , and free-form results are demonstrated in Appendix Table 5. We observe that most models struggle with reasoning MCQs compared to base and hypothetical questions. While some models perform poorly on reasoning and hypothetical MCQs, they tend to generate stronger responses for the same type of Freeform Questions, indicating improved performance when reasoning is emphasized over factual recall similar to observations of Chain of Thought experiments (Wei et al., 2022), as shown in Figure 5.
@@ 5.2 Results
We examine the impact of providing relevant context through RAG when answering questions. We explore two retrieval scenarios: one where the model retrieves from 13 books that were not used to generate the questions (rag-held-out), and another where it retrieves from the 5 books that the questions were derived from (rag-source). Retrieval from source textbooks consistently enhances performance across all tasks. For Gemini 1.5, we identify the book corresponding to the most relevant retrieved chunk under both the source and held-out settings, and include it as in-context information to answer the question. Additional context from source books yields a slight improvement over the no-context baseline. However, in both RAG and long-context scenarios, incorporating held-out books usually reduces performance, likely due to irrelevant or distracting content.
@@
Continued pertaining (cp-held-out) on the graduate textbooks leads to improved performance in both MCQs and Cloze question-answering tasks. Additionally, fine-tuning (ft-silver) on the ClimaQA-Silver dataset further enhances performance, often producing the best results after RAG on the source textbooks (rag-source) in most scenarios. Furthermore, Few-shot prompting yields marginal improvements in most cases.
@@
Finally, We observe that the BLEU and BERTScore metrics are slightly biased towards the model that was used for QA-generation (gpt-3.5-turbo) while this is not seen in the proposed Factual Accuracy metric 5. Overall, GPT-4o dominates across tasks, demonstrating superior performance compared to other models in this evaluation set.
@@ 6 Conclusion
The ClimaQA benchmark offers a comprehensive framework for evaluating language models in climate question-answering, addressing critical aspects such as reasoning, factual accuracy, and understanding of scientific terminology. By incorporating freeform, multiple-choice, and cloze task forms with different levels of complexity, the benchmark rigorously tests models across different dimensions of scientific inquiry. Furthermore, the use of advanced metrics, such as factual accuracy for freeform tasks and phrase similarity for cloze tasks, can provide a more nuanced assessment of model performance.
@@
Reproducibility Statement
The ClimaQA dataset, both gold and silver, is publicly available at Hugging Face [^1] https://huggingface.co/datasets/UCSD-GENIE/ClimaQA. While we cannot release the scraped textbook data due to copyright restrictions, the references to all textbooks used are provided in the appendixA.1, allowing for reconstruction of this dataset. Our complete codebase, including data generation pipeline, web-UI and model evaluation scripts, is available in our GitHub repository [^2] https://github.com/Rose-STL-Lab/genie-climaqa. The ClimaGen framework is fully reproducible; however, training the evaluator models requires domain expert inputs, which may introduce variability. Furthermore, reproduction of all parts of our framework requires appropriate API keys for external services.
@@ Appendix A Appendix
A.1 Textbook Dataset
Table 4: List of climate textbooks used in this work along with the number of pages and the number of context chunks. The extracted data was preprocessed to omit figures and tables. The content of each page was then split into overlapping context chunks of around 2000 characters.
Textbook Pages Chunks
Aerosol Measurement (Baron & Willeke, 2001) 937 1813
Aerosols and Climate (Carslaw, 2022) 854 1766
Airborne CCN Measurements (Trembath, 2013) 268 333
An Introduction to Clouds (Lohmann et al., 2016) 377 606
Atmospheric Chemistry and Physics (Seinfeld & Pandis, 2016) 1127 1789
Atmospheric Science (Wallace & Hobbs, 2006) 495 1085
Calculus of Variations (Gelfand & Fomin, 2012) 239 266
Clouds in the Perturbed Climate System (Heintzenberg & Charlson, 2009) 598 1060
Eloquent Science (Schultz, 2013) 420 725
Filtering Complex Turbulent Systems (Majda & Harlim, 2012) 358 475
Fundamental of Atmospheric Modeling (Jacobson, 2005) 826 1204
Geostatistics for Environmental Scientists (Webster & Oliver, 2007) 327 457
Global Physical Climatology (Hartmann, 2015) 481 715
Principles Of Planetary Climate (Pierrehumbert, 2010) 468 1096
Forests and Climate Change (De Wasseige et al., 2015) 119 250
Simulating Nature (Petersen, 2012) 205 386
Statistical Methods in the Atmospheric Sciences (Wilks, 2019) 806 1597
Stochastic Climate Models (Imkeller & von Storch, 2012) 411 594
To assess the proficiency of LLMs in climate science, we utilized graduate-level climate science textbooks as a reliable source of specialized knowledge. These textbooks were selected for their accurate and comprehensive representation of the technical terminology and nuanced theories integral to the field. The collection was curated from the virtual bookshelf of a professor in atmospheric physics and broadly represents a mixture of graduate and expert textbooks on the physical climate, with a particular focus on the role of aerosol in the climate system (which provides one of the key uncertainties in climate projections). The complete list of textbooks is provided in Table 4.
Figure 6: The distribution of ClimaQA-Gold questions over the corresponding source.
The five textbooks selected for question generation were carefully chosen to reflect different levels of breadth and depth (both technically and more qualitatively) across a range of important topics in climate science. These include Atmospheric Science, a comprehensive graduate-level textbook on atmospheric science; Aerosols and Climate, a brand-new and more detailed textbook on the role of aerosol in the climate system; An Introduction to Clouds, another relatively new textbook providing the latest research on the important role of clouds in the climate system; Geostatistics for Environmental Scientists, a more technical graduate-level textbook on geostatistics for environmental sciences; as well as Calculus of Variations, a classic calculus textbook to test the extent of the technical ability of the models. The exact distribution of questions from these textbooks is shown in Figure 6. The remaining textbooks were used in the rag-held-out experiments to test the models without directly exposing them to the sources of the benchmark.
A.2 More Experiments
A.2.1 Freeform Evaluation Results
Table 5: Performance analysis of various state-of-the-art LLMs on Freeform QA. Bold marks the max within a model’s variants and green highlights the overall column max.
@@ B R H O B R H O B R H O
gemma-27b-rag-source 0.462 0.573 0.552 0.529 0.902 0.911 0.906 0.906 0.84 0.85 0.85 0.85
gemma-27b-rag-held-out 0.349 0.496 0.524 0.456 0.867 0.889 0.896 0.885 0.58 0.73 0.77 0.69
@@
gpt-3.5-turbo-rag-source 0.528 0.601 0.561 0.563 0.908 0.919 0.921 0.916 0.85 0.86 0.87 0.86
gpt-3.5-turbo-rag-held-out 0.425 0.477 0.543 0.482 0.877 0.889 0.904 0.891 0.46 0.64 0.81 0.64
@@
llama3-70b-rag-source 0.441 0.574 0.604 0.540 0.909 0.916 0.917 0.914 0.82 0.81 0.84 0.82
llama3-70b-rag-held-out 0.341 0.475 0.519 0.445 0.872 0.890 0.902 0.888 0.45 0.62 0.79 0.62
@@
mixtral-8x22b-rag-source 0.458 0.598 0.610 0.555 0.905 0.916 0.912 0.911 0.78 0.82 0.86 0.82
mixtral-8x22b-rag-held-out 0.341 0.488 0.546 0.459 0.870 0.890 0.898 0.886 0.50 0.62 0.77 0.63
@@
llama3.1-8b-rag-source 0.509 0.550 0.534 0.531 0.905 0.920 0.911 0.912 0.77 0.80 0.83 0.80
llama3.1-8b-rag-held-out 0.392 0.469 0.507 0.456 0.873 0.893 0.901 0.889 0.52 0.65 0.72 0.63
llama3.1-8b-cp-held-out 0.420 0.455 0.496 0.457 0.876 0.892 0.902 0.890 0.53 0.70 0.71 0.65
llama3.1-8b-ft-silver 0.436 0.469 0.533 0.480 0.882 0.896 0.904 0.894 0.53 0.59 0.77 0.63
@@
mistral-7b-rag-source 0.541 0.455 0.448 0.482 0.898 0.912 0.907 0.906 0.77 0.82 0.84 0.81
mistral-7b-rag-held-out 0.413 0.356 0.399 0.390 0.868 0.884 0.896 0.883 0.54 0.60 0.73 0.62
mistral-7b-cp-held-out 0.295 0.176 0.189 0.221 0.880 0.890 0.899 0.890 0.52 0.70 0.70 0.64
mistral-7b-ft-silver 0.520 0.386 0.434 0.447 0.892 0.900 0.906 0.899 0.40 0.61 0.82 0.61
A.2.2 ClimaQA-Silver Experiments
In this section, we assess the models in their default settings on a subset of the ClimaQA-Silver dataset, consisting of 200 questions for each task type. The column-wise trends observed in the results are largely consistent with those from the ClimaQA-Gold dataset. However, a notable difference lies in the relative difficulty of the complexities in the MCQ task, even though the overall scores do not vary significantly. This discrepancy highlights the intricate nature of the MCQ generation process. Additionally, this variation may also be attributed to the relatively small number of questions in each column. Future work should focus on scaling up the expert validation process to enhance the quality of the automated annotation pipeline, thereby addressing these challenges and improving overall dataset reliability.
Table 6: Performance analysis of various LLMs on MCQs and Cloze QA in ClimaQA-Silver
Model MCQ Cloze
Base Reason Hypo Overall EM PS
gemma-27b 78.00 84.91 76.6 79.5 50.00 0.85
gpt-3.5-turbo 76.00 75.47 72.34 75.0 38.00 0.78
gpt-4o 88.00 84.91 78.72 85.00 60.50 0.88
llama3-70b 85.00 79.25 65.96 79.00 44.00 0.83
mixtral-8x22b 80.00 81.13 82.98 81.00 33.00 0.67
Table 7: Performance analysis of various LLMs on Freeform QA in ClimaQA-Silver
Model BLEU BERTScore Factual Acccuracy
B R H O B R H O B R H O
gemma-27b 0.392 0.441 0.365 0.398 0.870 0.880 0.886 0.876 0.78 0.86 0.85 0.81
gpt-3.5-turbo 0.467 0.523 0.545 0.500 0.885 0.892 0.907 0.892 0.71 0.79 0.84 0.76
gpt-4o 0.440 0.493 0.491 0.465 0.880 0.887 0.902 0.887 0.80 0.88 0.86 0.84
llama3-70b 0.335 0.474 0.569 0.425 0.874 0.888 0.904 0.885 0.77 0.82 0.75 0.78
mixtral-8x22b 0.394 0.516 0.485 0.447 0.877 0.890 0.904 0.887 0.80 0.88 0.83 0.82
A.2.3 Case Study
In this section, we present sample questions from the ClimaQA-Gold dataset along with responses from various LLMs and their corresponding evaluation scores. Freeform QA
In this example, the reference answer discusses increased global emissions in the absence of air quality measures. While the outputs of most LLMs align with this reference, the response from gemma-27b suggests a decline in global emissions. Note that both the Factual Accuracy and BLEU metrics penalize this response, whereas the BERTScore does not.
{mdframed}
[linewidth=1pt] Quesition: How might the global emission estimates have been different if air quality measures were not introduced mainly targeted at large, coal-burning installations before 2000?
Reference Answer: Without the introduction of air quality measures mainly targeted at large, coal-burning installations before 2000, the global emission estimates, particularly for SO2, may not have shown a decrease before 2000. The emissions could have continued to rise or remained at higher levels due to the ongoing industrial growth and lack of regulations to curb pollution from such sources.
gpt-3.5-turbo: Global emission estimates would likely have been higher if air quality measures targeting large, coal-burning installations were not introduced before 2000, as these measures have helped to reduce emissions of pollutants such as sulfur dioxide and particulate matter.
Bleu: 0.485, Bert: 0.901, Factual Accuracy: 0.988
gpt-4o: Without air quality measures targeting large, coal-burning installations before 2000, global emissions, particularly sulfur dioxide and particulate matter, would likely have been significantly higher, exacerbating air pollution and related health issues. This could have also contributed to more severe acid rain and potentially altered climate patterns due to increased atmospheric aerosols.
Bleu: 0.589, Bert: 0.898, Factual Accuracy: 0.982
llama3-70b: If air quality measures were not introduced before 2000, global emission estimates would likely be higher, as coal-burning installations would have continued to emit large amounts of pollutants, including particulate matter, sulfur dioxide, and nitrogen oxides, contributing to increased greenhouse gas emissions. This could have resulted in an additional 1-2 gigatons of CO2 equivalent emissions per year, exacerbating climate change impacts.
Bleu: 0.542, Bert: 0.888, Factual Accuracy: 0.968
mixtral-8x22b: If air quality measures hadn’t been introduced for large, coal-burning installations before 2000, global emission estimates would likely have been significantly higher due to unregulated emissions from these sources. This could have potentially accelerated global warming and worsened air quality on a global scale.
Bleu: 0.546, Bert: 0.910, Factual Accuracy: 0.988
gemma-27b: Global emission estimates would likely be significantly higher without those air quality measures. Reduced emissions from coal-burning power plants would have contributed less to the overall decline in global emissions observed since 2000.
Bleu: 0.364, Bert: 0.885, Factual Accuracy: 0.741
In the following example, the reference answer indicates that the distance from the sun influences surface pressure, with higher pressures during summer. Responses from gpt-3.5-turbo and mixtral-8x22b contradict this statement but still achieve high BLEU and BERTScore values. The Factual Accuracy metric, however, correctly identifies and penalizes this inconsistency. Notably, gpt-4o provides a correct answer yet receives a low BLEU score, as BLEU is not robust against paraphrasing.
{mdframed}
[linewidth=1pt] Quesition: How does the distance from the sun affect the surface pressure on Mars?
Reference Answer: The surface pressure on Mars varies seasonally from 7.0 hPa during the austral winter, when Mars is farthest from the sun, to 9.0 hPa during the austral summer.
gpt-3.5-turbo: The distance from the sun does not directly affect the surface pressure on Mars; instead, the surface pressure on Mars is primarily influenced by the planet’s atmospheric composition and temperature.
Bleu: 0.388, Bert: 0.880, Factual Accuracy: 0.175
gpt-4o: The distance from the Sun affects Mars’ temperature, which in turn influences the sublimation of its polar ice caps, impacting atmospheric pressure. During warmer periods, more CO2 sublimates, increasing surface pressure, while colder periods result in lower pressure.
Bleu: 0.253, Bert: 0.865, Factual Accuracy: 0.980
llama3-70b: The distance from the Sun affects the surface pressure on Mars by controlling the amount of atmospheric gas that freezes out or escapes to space, with more gas freezing out at greater distances. As a result, Mars’ surface pressure is about 1% of Earth’s, due to its average distance from the Sun being about 1.5 times greater than Earth’s.
Bleu: 0.262, Bert: 0.861, Factual Accuracy: 0.896
mixtral-8x22b: The distance from the sun does not significantly affect the surface pressure on Mars. Instead, Mars’ thin atmosphere and low gravity are the primary factors influencing its surface pressure.
Bleu: 0.403, Bert: 0.874, Factual Accuracy: 0.051
gemma-27b: Mars is farther from the sun than Earth, receiving less solar radiation. This weaker solar heating results in a much lower surface pressure on Mars compared to Earth.
Bleu: 0.458, Bert: 0.870, Factual Accuracy: 0.905
Interestingly, gemma-27b provides an irrelevant response that nevertheless receives high scores. While the response is factually accurate, it fails to address the question, illustrating limitations in existing evaluation metrics. This highlights the need for more fine-grained evaluation methodologies, which future research should aim to develop.
Cloze QA
In the following examples, the models generate answers that differ from the reference answer but remain correct within the given context. This is effectively captured by the phrase similarity metric, which assigns higher scores to responses that are contextually aligned with the reference answer.
{mdframed}
[linewidth=1pt] Question: Aqueous-phase oxidation of $\langle mask\rangle$ to form sulfuric acid in cloud droplets is a crucial process that contributes significantly to global sulfate mass and has long been recognized as important for aerosols and climate.
Reference Answer: SO2
gpt-3.5-turbo: Sulfur, Phrase Similarity: 0.819
gpt-4o: sulfur dioxide, Phrase Similarity: 0.885
llama3-70b: Sulfur, Phrase Similarity: 0.819
mixtral-8x22b: sulfur dioxide, Phrase Similarity: 0.885
gemma-27b: sulfur dioxide, Phrase Similarity: 0.885 {mdframed}[linewidth=1pt] Question: The rate of change in the air parcel’s moist static energy due to $\langle mask\rangle$ is determined by the difference between the moist static energies of the environment and the cloud air parcel.
Reference Answer: entrainment
gpt-3.5-turbo: evaporation, Phrase Similarity: 0.681
gpt-4o: entrainment, Phrase Similarity: 1.000
llama3-70b: entrainment, Phrase Similarity: 1.000
mixtral-8x22b: mixing, Phrase Similarity: 0.726
gemma-27b: entrainment, Phrase Similarity: 1.000
A.3 Training Details